When 1 mole of Zn metal is added to 2M HCl solution at 25 °C and 100 kPa pressure, the metal dissolves and hydrogen gas is evolved. At the same time, 152.3 kJ of heat energy are released. a. Calculate ∆V for the reaction
1
Expert's answer
2020-09-05T14:42:08-0400

When 1 mole of Zn metal is added to 2M HCl solution .


Then 1 mol of H2 gas is formed due to this gas work done prformed .


energy = 152.3 * 1000 = ( 100 * 1000) * delta V .


delta V = 152300 / 100000 = 1.523 .